OpenStudy (anonymous):

what is the derivative of tan^3(x)

OpenStudy (amistre64):

looks like power, chain, trig to me

OpenStudy (campbell_st):

well make a little substitution u = tan (x) then y = u^3 find dy/du and du/dx then dy/dx = dy/du * du/dx

OpenStudy (anonymous):

im still lost

OpenStudy (amistre64):

im assuming you have a rudimentary knowledge of a derivative. it might be easier to view this as: \[[trig(x)]^n\] power rule it \[n~[trig(x)]^{n-1}\] but a chain pops out a derivative of the insides \[n~[trig(x)]^{n-1}~~[trig(x)]'\] so it amounts to determining the derivative of a trig now